# Tight PAC-Bayesian Risk Certificates for Contrastive Learning

### v2

@@ 3.2 Bound on the Downstream Classification Loss
We apply this lemma for $N=m-1$ and $L=\frac{1}{\sqrt{\tau}}$. For convenience, we use $K=m-1$ to denote the number of negative samples. We obtain
$\mathbb{E}_{\begin{subarray}{c}\{y^{\prime}\}\end{subarray}}\left[\operatorname{LSE}(\mathbf{z}^{\mu})\right]$ $\overset{(a)}{\geq}-\mathbb{E}_{\begin{subarray}{c}\{y^{\prime}\}\end{subarray}}\left[\operatorname{LSE}(-\mathbf{z}^{\mu})\right]+2\log(K)$
$\overset{(b)}{\geq}-\log\sum_{\{y^{\prime}\}}\mathbb{E}_{\begin{subarray}{c}y^{\prime}\end{subarray}}\left[\exp(-f(x)^{\top}\mu_{y^{\prime}})\right]+2\log(K)$
=$\overset{(c)}{=}-\log\sum_{c\in\mathcal{C}}\exp(-f(x)^{\top}\mu_{c})\pi(c)+2\log(K)$
=$\overset{(d)}{=}-\log\frac{K}{C}\sum_{c\in\mathcal{C}}\exp(-f(x)^{\top}\mu_{c})+2\log(K)$
=$\overset{(e)}{=}-\operatorname{LSE}\left(\left\{-f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}\right)+\log(K)+\log(C)$
$\overset{(f)}{\geq}\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}\right)-2\log\left(C\cosh\frac{1}{\tau}\right)+\log(K)+\log(C)$
=$\overset{(g)}{=}\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}\right)-\Delta$
where (a) directly comes from lemma 10, (b) uses Jensen’s inequality, (c) uses the distribution $\pi(c)$ over the classes and $\mathcal{C}$ represents the set of classes, (d) assumes that $\pi(c)=\frac{1}{C}$ where $C$ corresponds to the number of classes, (e) uses the definition of the $\operatorname{LSE}$ function, (f) is another application of lemma 10, and (g) uses the definition $\Delta:=\log\left(\frac{C}{K}\operatorname{cosh}^{2}(\frac{1}{\tau})\right)$.
@@
$L(f)$ ${\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}\right)-\Delta$
@@
Taking the minimum over all linear classifiers finishes the proof:
$\min_{W\in\mathbb{R}^{C\times d}}L_{\mathrm{CE}}(f,W)\leq L(f)+\frac{\sigma}{\tau}+\Delta.$
The second part of the theorem is an extension of lemma 9 obtained by removing the temperature scaling from the $\operatorname{LSE}$ term. The result is stated in the following lemma.
@@
$\min_{W\in\mathbb{R}^{C\times d}}L_{\mathrm{CE}}(f,W)\leq\sigma+\tau L(f)+\Delta^{\prime},$
where $\Delta^{\prime}=\log(C)+\min\left\{\log(\operatorname{cosh}^{2}(1)),\tau\log\left(\frac{C}{K}\operatorname{cosh}^{2}\left(\frac{1}{\tau}\right)\right)\right\}$.
Proof.
We begin by deriving few useful inequalities that intuitively show that the $\operatorname{LSE}$ function acts as a smooth maximum.
For $N\in\mathbb{N}^{*}$, $\mathbf{x}\in\mathbb{R}^{N}$ with $\mathbf{x}=(x_{1},\ldots,x_{N})$,
and for any $t>0$,
@@
The inequality is verified by observing that if $m=\max_{i}x_{i}$, then we have the following inequality $\exp(tm)\leq\sum_{i=1}^{N}\exp\left(tx_{i}\right)\leq N\exp(tm)$. Applying the logarithm to the inequality gives the first result.
Define $W^{\mu}:=[\mu_{1}\cdots\mu_{C}]^{\top}$. The bound presented in lemma 11 follows from two applications of (11) at different steps of the proof of lemma 9. On one hand, applying (11) to the term $\operatorname{LSE}(\mathbf{z})$ introduces a factor of $\log(K)$ With $\mathbf{u}:=\left\{f(x)^{\top}f(x^{\prime})\right\}_{x^{\prime}\in X}$, this yields $\operatorname{LSE}(\mathbf{z})\geq\frac{1}{\tau}\left(\operatorname{LSE}(\mathbf{u})-\log(K)\right)$. Applying lemma 9 to $\operatorname{LSE}(\mathbf{u})$ for $L=1$ gives $\operatorname{LSE}(\mathbf{u})\geq\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}\right\}_{c\in[C]}\right)-\log\left(\frac{C}{K}\operatorname{cosh}^{2}(1)\right)$. Combining this with the previous bound, we derive
$L_{\mathrm{CE}}(f,W^{\mu})\leq\tau L(f)+\sigma+\log(C\operatorname{cosh}^{2}(1)).$ (12)
On the other hand, applying (11) to the term $\operatorname{LSE}\left(\mathbf{z}^{\mu}\right)$ introduces a factor of $\log(C)$, leading to
$\tau\operatorname{LSE}\left(\mathbf{z}^{\mu}\right)\geq\operatorname{LSE}\left(\mathbf{u}^{\mu}\right)-\log(C).$ Plugging this into the proof of lemma 9, we obtain
$L_{\mathrm{CE}}(f,W^{\mu})\leq\tau L(f)+\sigma+\tau\Delta+\log(C).$ (13)
Finally, taking the minimum over both refined bounds eq. 12 and eq. 13 and then over all linear classifiers completes the proof.
Combining lemma 9 and lemma 11, as well as, rewriting with the following notation
$\beta(f,\sigma):=\frac{\sigma}{\tau}+L(f)+\Delta$ completes the proof of theorem 5.
@@ 3.3 Risk Certificate for Contrastive Zero-One Risk
In the previous sections, we demonstrated that non-vacuous risk certificates can be obtained for the SimCLR loss and that the SimCLR loss acts as a surrogate loss for downstream classification. Similarly to Nozawa et al. [32], we extend our analysis to the contrastive zero-one risk. Recall the population contrastive zero-one risk $R(f)$ in (5). Let $\widehat{R}_{S}(f)$ denote the empirical counterpart of $R(f)$. We extend both risk certificates from section 3.1 to the contrastive zero-one risk. First, we present theorem 6 corresponding to the extension of theorem 3.
Theorem 6 (Extension of theorem 3).
@@
For the first part, we follow the proof from lemma 1, except for: we observe that $\delta_{i}(x_{i})\leq 1$ and $\delta_{j}(x_{i})\leq\frac{1}{m-1}$, and thus we obtain $\sup_{x_{1},\ldots,x_{n},x_{i}^{\prime}\in X}\left|\Delta\phi(x_{i})\right|\leq\frac{2}{n}$. For the PAC-Bayes bound, one can directly apply theorem 3 with $C=2$. ∎
One can also extend theorem 4, as stated below.
Theorem 7 (Extension of theorem 4).
@@
where $\gamma=\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{2(m-1)\alpha}}$.
@@
We follow the proof from theorem 4: (1) the sum containing the negative samples becomes $S(x,X^{-}):=\sum_{x^{\prime}\in X^{-}}\mathbb{I}_{\left\{f(x)^{\top}f(x^{+})<f(x)^{\top}f(x^{\prime})\right\}}$ (2) we obtain $\varepsilon=\sqrt{\frac{(m-1)\log\left(\frac{1}{\delta_{c}}\right)}{2}}$ since $\mathbb{I}_{\{\cdot\}}\leq 1$ and we have $B_{\ell}=1$ (3) the upper bound on the contrastive zero-one population risk by an intermediate loss becomes $R(Q)\leq\widetilde{R}(Q)+\frac{\varepsilon}{m-1}+\delta_{c}$ and the PAC-Bayes bound is computed using $B=1$ (4) Combining everything, setting $\gamma=\frac{\varepsilon}{m-1}$ and taking the infimum over $\alpha$ completes the proof. ∎
@@ 4 Experiments
In this section, we describe the experimental setup and empirically demonstrate that our risk certificates improved upon previous risk certificates through experiments on the CIFAR-10 dataset. For a comprehensive overview of all applicable previous bounds, please refer to the supplementary materials. The code for our experiments is publicly available  [^1] Code available at https://github.com/anna-vanelst/simclr-pb. Experimental results on the MNIST dataset and additional experimental details are available in the supplementary materials.
@@
We use two popular benchmarks: (1) CIFAR-10, which consists of 50,000 training images and 10,000 test images [23], and (2) MNIST, which consists of 60,000 training images and 10,000 test images [25], as provided in torchvision [27]. The images are preprocessed by normalizing all pixels per channel based on the training data. Data augmentation includes random cropping, resizing (with random flipping), and color distortions, as detailed in Appendix H [7]. We employ a 7-layer convolutional neural network (CNN) with max-pooling every two layers for CIFAR-10 experiments and a 3-layer CNN for MNIST experiments. We use a 2-layer MLP projection head to project to a 128-dimensional latent space, with a feature dimensionality of 2048 for CIFAR-10 and 512 for MNIST. ReLU activations are used in each hidden layer. The mean parameters $\mu_{0}$ of the prior are initialized randomly from a truncated centered Gaussian distribution with a standard deviation of $1/\sqrt{n_{\text{in}}}$, where $n_{\text{in}}$ is the dimension of the inputs to a particular layer, truncating at ±$\pm 2$ standard deviations [34]. The prior distribution scale (standard deviation $\sigma_{0}$) is selected from $\{0.01,0.05,0.1\}$.
@@
The learning and certification strategy involves three steps: (1) choose or learn a prior from a subset of the dataset; (2) learn a posterior on the entire training dataset; (3) evaluate the risk certificate for the posterior on a subset of the dataset independent of the prior. We experiment with two types of priors: informed and random. The informed prior is learned using a subset of the training dataset via empirical risk minimization or PAC-Bayes objective minimization. The posterior is initialized to the prior and learned using the entire training dataset by PAC-Bayes objective minimization. We use the PAC-Bayes with Backprop (PBB) procedure [35] and the following $f_{\text{classic}}$ objective : $f_{\text{classic }}(Q)=\frac{1}{B}\widehat{L}_{S}(Q)+\sqrt{\frac{\eta\mathrm{KL}\left(Q\|P\right)+\log\left(\frac{\sqrt{n}}{\delta}\right)}{2n}}$, where $\eta$ in $[0,1]$ is a coefficient introduced to control the influence of the KL term in the training objective, called the KL penalty. We use a KL penalty term of $10^{-6}$ for learning the prior and no penalty term for learning the posterior. We use SGD with momentum as optimizer and we perform a grid search for momentum values in $\{0.8,0.85,0.90,0.95\}$ and learning rates in $\{0.1,0.5,1.0,1.5\}$. Training was conducted for 100 epochs, and we selected the hyperparameters that give the best risk certificates. We experiment with different temperatures selected $\{0.2,0.5,0.7,1\}$. Unless otherwise specified, experiments are run using a probabilistic prior with the simplified SimCLR loss, a batch size of $m=250$, and 80% of the data for training the prior.
@@
Since $\widehat{L}_{S}(Q)$ is intractable, the final risk certificates are computed using Monte Carlo weight sampling. Specifically, we approximate $\widehat{L}_{S}(Q)$ using the empirical measure $\widehat{Q}_{p}=\sum_{j=1}^{p}\delta_{W_{j}}$, where $W_{1},\ldots,W_{p}\sim Q$ are i.i.d. samples. We compute all risk certificates with $\delta=0.04$, and $p=100$ Monte Carlo model samples. We report the risk certificates for both the contrastive loss and the contrastive zero-one risk using theorem 3 and theorem 4 for the contrastive loss, and theorem 6 and theorem 7 for the zero-one risk. To find the best value of $\alpha$ for theorem 4 and theorem 7, we perform a grid search over $\{0.1,0.2,0.3,0.4,0.5\}$. We observe that $\alpha=0.4$ provides the tightest risk certificates. Our risk certificates are compared with existing ones, including the PAC-Bayes-kl, classic PAC-Bayes or Catoni’s bounds over i.i.d. batches and the $f$-divergence PAC-Bayes bound, as detailed in the supplementary materials.
@@
We assess the quality of the learned representations through linear evaluation [7]: we report the cross-entropy loss and top-1 accuracy of linear classifiers trained on features either before or after the projection head. The classifiers are trained for $20$ epochs on the image classification task ($C=10$) using the Adam optimizer with a learning rate of $0.01$. Finally, we report the bounds on the downstream classification loss derived from theorem 5 and compare it with the state of the art [4].
Risk Certificate SimCLR Loss Contrastive 0-1 Risk
@@
Table 2: Comparison of risk certificates for the SimCLR loss and contrastive zero-one risk using different PAC-Bayes bounds for varying temperature values on CIFAR-10. kl bound (iid) refers to the standard PAC-Bayes-kl bound computed over i.i.d. batches, Catoni’s bound (iid) refers to Catoni’s PAC-Bayes bound computed over i.i.d. batches [32], Classic bound (iid) refers to the classic PAC-Bayes bound computed over i.i.d. batches, and Nozawa et al. refers to the PAC-Bayes bound based on $f$-divergence. Although Nozawa et al.’s bound uses $\chi^{2}$ divergence, we use $\operatorname{KL}$ divergence, which already results in vacuous bounds and would not improve with $\chi^{2}$, since $\operatorname{KL}(P\|Q)\leq\chi^{2}(P\|Q)$. We report test losses and observe that our bounds are remarkably tight. We also report the complexity term, $\operatorname{KL}/n$, where $\operatorname{KL}$ represents the Kullback-Leibler divergence between the prior and posterior distributions, and $n$ is the dataset size used to compute the risk certificate.
Proj. $\tau=1$ $\tau=0.7$ $\tau=0.5$ $\tau=0.2$
@@
Table 3: Comparison of upper bounds on downstream classification loss on CIFAR-10. We compare the original bound from Bao et al. with our refined bound (theorem 5). The supervised loss of the linear classifier trained on the projected features is reported, as it is directly related to the theoretical upper bound. Additionally, we report the supervised loss of the linear classifier trained on the features after removing the projection head. We empirically observe that the supervised loss of a linear classifier trained on the full features (without projection) is consistently lower than the loss of a linear classifier trained on the projected features, aligning with previous findings [7]. For reference, we also include top-1 accuracy.

### v4

@@ 3.2 Bound on the Downstream Classification Loss
We apply this lemma for $N=K$ and $L={1/\sqrt{\tau}}$. We obtain
$\mathbb{E}_{Y}\left[\operatorname{LSE}(\mathbf{z}^{\mu})\right]$ $\overset{(a)}{\geq}-\mathbb{E}_{Y}\left[\operatorname{LSE}(-\mathbf{z}^{\mu})\right]+2\log(K)$ (16)
$\overset{(b)}{\geq}-\log\sum_{y^{\prime}\in Y}\mathbb{E}_{\begin{subarray}{c}y^{\prime}\end{subarray}}\left[\exp(-f(x)^{\top}\mu_{y^{\prime}}/\tau)\right]+2\log(K)$ (17)
$\overset{(c)}{=}-\log K\sum_{c\in\mathcal{C}}\exp(-f(x)^{\top}\mu_{c}/\tau)\pi(c)+2\log(K)$ (18)
$\overset{(d)}{\geq}-\log{K\pi^{*}}\sum_{c\in\mathcal{C}}\exp(-f(x)^{\top}\mu_{c}/\tau)+2\log(K)$ (19)
$\overset{(e)}{=}-\operatorname{LSE}\left(\left\{-f(x)^{\top}\mu_{c}/\tau\right\}_{c\in\mathcal{C}}\right)+\log(K)-\log(\pi^{*})$ (20)
$\overset{(f)}{\geq}\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}/\tau\right\}_{c\in\mathcal{C}}\right)-2\log\left(C\cosh\frac{1}{\tau}\right)+\log(K)-\log(\pi^{*})$ (21)
$\overset{(g)}{=}\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}/\tau\right\}_{c\in\mathcal{C}}\right)-\Delta\,$, (22)
where (a) directly comes from lemma˜10, (b) applies Jensen’s inequality to the convex function $x\mapsto-\log x$ and then the linearity of expectation, (c) uses the distribution $\pi(c)$ over the classes and $\mathcal{C}$ represents the set of classes, (d) uses the bound $\pi(c)\leq\pi^{*}$ for all $c\in\mathcal{C}$ and the monotonicity of $\log$, (e) uses the definition of the $\operatorname{LSE}$ function, (f) is another application of lemma˜10, and (g) uses the definition $\Delta:=\log\left(\frac{C^{2}\pi^{*}}{K}\operatorname{cosh}^{2}(\frac{1}{\tau})\right)$.
@@
$L(f)$ ${\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\operatorname{LSE}\left(\left\{\frac{f(x)^{\top}\mu_{c}}{\tau}\right\}_{c\in\mathcal{C}}\right)-\Delta$
@@
Taking the minimum over all linear classifiers finishes the proof. ∎
The second part of the theorem is an extension of lemma˜9 obtained by removing the temperature scaling from the $\operatorname{LSE}$ term. The result is stated in the following lemma.
@@
$\min_{W\in\mathbb{R}^{C\times d}}L_{\mathrm{CE}}(f,W)\leq\sigma+\tau L(f)+\tau\Delta+\alpha,$
where $\alpha=\log(C)+\min\left\{0,\log(C\pi^{*}\operatorname{cosh}^{2}(1))-\tau\Delta\right\}$.
Sketch of Proof.
For a detailed proof of the lemma, we refer the reader to the supplementary materials. We first derive a useful inequality that show that the $\operatorname{LSE}$ function acts as a smooth maximum. For $N\in\mathbb{N}^{*}$, $\mathbf{x}\in\mathbb{R}^{N}$ with $\mathbf{x}=(x_{1},\ldots,x_{N})$, and for any $t>0$,
@@
The inequality is verified by observing that if $m=\max_{i}x_{i}$, then we have the following inequality $\exp(tm)\leq\sum_{i=1}^{N}\exp\left(tx_{i}\right)\leq N\exp(tm)$. Applying the logarithm gives the result.
Define $W^{\mu}:=[\mu_{1}\cdots\mu_{C}]^{\top}$. The bound presented in lemma˜11 follows from two applications of (23) at different steps of the proof of lemma˜9. On one hand, applying (23) to the term $\operatorname{LSE}(\mathbf{z})$ introduces a factor of $\log(K)$ . With $\mathbf{u}:=\left\{f(x)^{\top}f(x^{\prime})\right\}_{x^{\prime}\in X}$, this yields $\operatorname{LSE}(\mathbf{z})\geq\frac{1}{\tau}\left(\operatorname{LSE}(\mathbf{u})-\log(K)\right)$. Following a proof similar to lemma˜9 with $L=1$, we obtain a lower bound on $\operatorname{LSE}(\mathbf{u})$: $\operatorname{LSE}(\mathbf{u})\geq\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}\right\}_{c\in[C]}\right)-\log\left(\frac{C^{2}\pi^{*}}{K}\operatorname{cosh}^{2}(1)\right)$. Combining this with the previous bound, we derive
$L_{\mathrm{CE}}(f,W^{\mu})\leq\sigma+\tau L(f)+\tau\Delta+\log(C)+\log(C\pi^{*}\operatorname{cosh}^{2}(1))-\tau\Delta.$ (24)
On the other hand, applying (23) to the term $\operatorname{LSE}(\mathbf{z}^{c})$ with $\mathbf{z}^{c}:=\left\{{f(x)^{\top}\mu_{c}/\tau}\right\}_{c\in\mathcal{C}}$ introduces a factor of $\log(C)$, leading to $\tau\operatorname{LSE}\left(\mathbf{z}^{c}\right)\geq\operatorname{LSE}\left(\mathbf{u}^{c}\right)-\log(C),$ where $\mathbf{u}^{c}:=\left\{f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}$. Plugging this into the proof of lemma˜9, we obtain
$L_{\mathrm{CE}}(f,W^{\mu})\leq\sigma+\tau L(f)+\tau\Delta+\log(C).$ (25)
Finally, taking the minimum over both refined bounds eq.˜24 and eq.˜25 and then over all linear classifiers completes the proof. ∎
Combining lemma˜9 and lemma˜11, as well as rewriting with the following notation $\beta(f,\sigma):={\sigma/\tau}+L(f)+\Delta$ completes the proof of theorem˜5.
@@ 3.3 Risk Certificate for Contrastive Zero-One Risk
In the previous sections, we demonstrated that non-vacuous risk certificates can be obtained for the SimCLR loss and that the SimCLR loss acts as a surrogate loss for downstream classification. Similarly to Nozawa et al. [36], we extend our analysis to the contrastive zero-one risk. Recall the population contrastive zero-one risk $R(f)$ in (5). Let $\widehat{R}_{S}(f)$ denote the empirical counterpart of $R(f)$. We extend both risk certificates from section˜3.1 to the contrastive zero-one risk. First, we present theorem˜6 corresponding to the extension of theorem˜3.
Theorem 6 (Extension of theorem˜3).
@@
For the first part, we follow the proof from lemma˜1, with the only difference that $\delta_{i}(x_{i})\leq 1$ and $\delta_{j}(x_{i})\leq\frac{1}{m-1}$, and thus we obtain $\sup_{x_{1},\ldots,x_{n},x_{i}^{\prime}\in X}\left|\Delta\phi(x_{i})\right|\leq\frac{2}{n}$. For the PAC-Bayes bound, we can follow the proof of theorem˜3 and use $C=2$. ∎
One can also extend theorem˜4, as stated below.
Theorem 7 (Extension of theorem˜4).
@@
where $\gamma=\sqrt{{(\log\left({2/\delta}\right))/2(m-1)\alpha}}$.
@@
We follow the proof from theorem˜4: (1) the sum containing the negative samples becomes $S(x,X^{-}):=\sum_{x^{\prime}\in X^{-}}\mathbb{I}_{\left\{f(x)^{\top}f(x^{+})<f(x)^{\top}f(x^{\prime})\right\}}$; (2) since $\mathbb{I}_{\{\cdot\}}\leq 1$, we obtain $\varepsilon=\smash{\sqrt{((m-1)\log({2/\delta}))/2}}$ and we have $B_{\ell}=1$; (3) the upper bound on the contrastive zero-one population risk by an intermediate loss becomes $R(Q)\leq\widetilde{R}(Q)+\frac{\varepsilon}{m-1}+\frac{\delta}{2}$ and the PAC-Bayes bound is computed using $B=1$; (4) we set $\gamma=\frac{\varepsilon}{m-1}$ and take the infimum over $\alpha$. ∎
@@ 4 Experiments
In this section, we describe the experimental setup and empirically demonstrate that our risk certificates improved upon previous risk certificates through experiments on the CIFAR-10 dataset. For a comprehensive overview of all applicable previous bounds, please refer to the supplementary materials. The code for our experiments is available in PyTorch [39]. Experimental results on the MNIST dataset and additional experimental details are available in the supplementary materials.
@@
We use two popular benchmarks: (1) CIFAR-10, which consists of 50,000 training images and 10,000 test images [27], and (2) MNIST, which consists of 60,000 training images and 10,000 test images [29], as provided in torchvision [31]. The images are preprocessed by normalizing all pixels per channel based on the training data. Data augmentation includes random cropping, resizing (with random flipping), and color distortions, as detailed in Appendix H [7]. We employ a 7-layer convolutional neural network (CNN) with max-pooling every two layers for CIFAR-10 experiments and a 3-layer CNN for MNIST experiments. We use a 2-layer MLP projection head to project to a 128-dimensional latent space, with a feature dimensionality of 2048 for CIFAR-10 and 512 for MNIST. ReLU activations are used in each hidden layer. The mean parameters $\mu_{0}$ of the prior are initialized randomly from a truncated centered Gaussian distribution with a standard deviation of $1/\sqrt{n_{\text{in}}}$, where $n_{\text{in}}$ is the dimension of the inputs to a particular layer, truncating at $\pm 2$ standard deviations [40]. The prior distribution scale (standard deviation $\sigma_{0}$) is selected from $\{0.01,0.05,0.1\}$.
@@
The learning and certification strategy involves three steps: (1) choose or learn a prior from a subset of the dataset; (2) learn a posterior on the entire training dataset; (3) evaluate the risk certificate for the posterior on a subset of the dataset independent of the prior. We experiment with two types of priors: informed and random. The informed prior is learned using a subset of the training dataset via empirical risk minimization or PAC-Bayes objective minimization. The posterior is initialized to the prior and learned using the entire training dataset by PAC-Bayes objective minimization. We use the PAC-Bayes with Backprop (PBB) procedure [41] and the following $f_{\text{classic}}$ objective : $f_{\text{classic }}(Q)=(1/B)\widehat{L}_{S}(Q)+\smash{\sqrt{(\eta\mathrm{KL}\left(Q\|P\right)+\log({\sqrt{n}/\delta}))/2n}}$, where $\eta$ in $[0,1]$ is a coefficient introduced to control the influence of the KL term in the training objective, called the KL penalty. We use a KL penalty term of $10^{-6}$ for learning the prior and no penalty term for learning the posterior. We use SGD with momentum as optimizer and we perform a grid search for momentum values in $\{0.8,0.85,0.90,0.95\}$ and learning rates in $\{0.1,0.5,1.0,1.5\}$. Training was conducted for 100 epochs, and we selected the hyperparameters that give the best risk certificates. We experiment with different temperatures selected $\{0.2,0.5,0.7,1\}$. Unless otherwise specified, experiments are run using a probabilistic prior with the simplified SimCLR loss, a batch size of $m=250$, and 80% of the data for training the prior.
@@
Since $\widehat{L}_{S}(Q)$ is intractable, the final risk certificates are computed using Monte Carlo weight sampling. Specifically, we approximate $\widehat{L}_{S}(Q)$ using the empirical measure $\widehat{Q}_{p}=\sum_{j=1}^{p}\delta_{W_{j}}$, where $W_{1},\ldots,W_{p}\sim Q$ are i.i.d. samples. We compute all risk certificates with $\delta=0.04$, and $p=100$ Monte Carlo model samples. We report the risk certificates for both the contrastive loss and the contrastive zero-one risk using theorem˜3 and theorem˜4 for the contrastive loss, and theorem˜6 and theorem˜7 for the zero-one risk. To find the best value of $\alpha$ for theorem˜4 and theorem˜7, we perform a grid search over $\{0.1,0.2,0.3,0.4,0.5\}$. We observe that $\alpha=0.4$ provides the tightest risk certificates. Our risk certificates are compared with existing ones, as detailed in table˜2 and the supplementary materials.
@@
We assess the quality of the learned representations through linear evaluation [7]: we report the cross-entropy loss and top-1 accuracy of linear classifiers trained on features either before or after the projection head. The classifiers are trained for $20$ epochs on the image classification task ($C=10$) using the Adam optimizer with a learning rate of $0.01$. Finally, we report the bounds on the downstream classification loss derived from theorem˜5 and compare it with the state of the art [3].
Risk Certificate
SimCLR Loss Contrastive 0-1 Risk
@@
Table 2: Comparison of risk certificates for the SimCLR loss and contrastive zero-one risk using different PAC-Bayes bounds for varying temperature values on CIFAR-10. kl bound (iid) refers to the standard PAC-Bayes-kl bound computed over i.i.d. batches (see section˜3.1), Catoni’s bound (iid) refers to Catoni’s PAC-Bayes bound computed over i.i.d. batches [36], Classic bound (iid) refers to the classic PAC-Bayes bound computed over i.i.d. batches, and Nozawa et al. refers to the PAC-Bayes bound based on $f$-divergence. Although Nozawa et al.’s bound uses $\chi^{2}$ divergence, we use $\operatorname{KL}$ divergence, which already results in vacuous bounds and would not improve with $\chi^{2}$, since $\operatorname{KL}(P\|Q)\leq\chi^{2}(P\|Q)$. We report test losses and observe that our bounds are remarkably tight. We also report the complexity term, $\operatorname{KL}/n$, where $\operatorname{KL}$ represents the Kullback-Leibler divergence between the prior and posterior distributions, and $n$ is the dataset size used to compute the risk certificate.
Proj.
$\tau=1$ $\tau=0.7$ $\tau=0.5$ $\tau=0.2$
@@
Table 3: Comparison of upper bounds on downstream classification loss on CIFAR-10. We compare the original bound from Bao et al. with our refined bound (theorem˜5). The supervised loss of the linear classifier trained on the projected features is reported, as it is directly related to the theoretical upper bound. Additionally, we report the supervised loss of the linear classifier trained on the features after removing the projection head. We empirically observe that the supervised loss of a linear classifier trained on the full features (without projection) is consistently lower than the loss of a linear classifier trained on the projected features, aligning with previous findings [7]. For reference, we also include top-1 accuracy.